OpenStudy (anonymous):

Ginny's class ordered 4 pizzas. They ate seven-twelfths of each pizza. How many pizzas did they eat? tell me

OpenStudy (anonymous):

7/12 x4=28/12

OpenStudy (anonymous):

7/12 x4=28/12 that is my answer yes or no.

OpenStudy (anonymous):

4

OpenStudy (radar):

Yes